Shift report 20020124 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 24 2002 - 00:06:58 EST

  • Next message: Christian Holm Christensen: "Re: FYI-will affect us..16/17 68/69"

    ------------------------------------------------------------
    Date: Thu Jan 24 00:06:57 2002, Shift: 16:00-24:00
    Supervisor: BHS, with: 
    Log book#12 pages 111 - 114
    
    Shift Summary:
    ------------------------------------------------------------
    So - the final shift report, the final official brahms
    shift for this run!
    
    The last evening has been a good one - we got a beautiful
    store at 16:15 which has lasted all through the shift. The
    final program set by Flemming and Dana was easily completed.
    
    The one problem we had was with D1 - it did not like being 
    run at full field (power overheated due to a problem with 
    the cooling...) and so the poor mag-guys spent a total
    of 2.5 hrs playing with the thing.
    
    Breakdown of data runs:
    
    All have MRS@90, FS@3 degrees
    
    Run    FS    MRS   Trig 3   Trig 4   Trig 6
    6750   1A    6B      8k       7k      30k  
    6752   1A    6B     10k       8k      32k  
    6754   1/2A  6B     16k      77k     130k
    6755   1/3A 12B      7k     114k     167k
    6756   1/3B 12B      7k     163k     222k
    6757   1/2B 12B     20k     272k     430k
    
    Runs with MRS only due to the D1 problems:
    6751   off   6B     10k       9k      39k
    6753   off   6B     49k     195k     382k 
    
    Magnets are now on standby and all anodes have been turned
    off/parked, but all the rest of the stuff is left on in case
    anyone wants to do some work with the pp2pp-specific beam
    tomorrow.
    Yeeeeeha :-)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Jan 24 2002 - 00:07:47 EST